Please help me solve this problem (2/3)(x+9)=x+5
This is from a worksheet I have in my college algebra class. I tried to solve it and this is what I come up with: 2/3x+9 2/3 =x+5
-2/3x -2/3
9 2/3=1/3x+5
-5 -5
4 2/3=1/3x x=14/3

Answer by checkley75(3666)

(2/3)(x+9)=x+5
2/3x+2/3*9=x+5
2/3x+18/3=x+5
2/3x-x=5-6
-1/3x=-1
x=-1/(-1/3)
x=-1*-3
x=3 answer
proof
(2/3)(3+9)=3+5
2/3*12=8
24/3=8
8=8